Suppose you have a 50.00ml of 0.0100 M chloroacetiPlease show steps and thought process.
Answer #1


weak acid pH = 1/2(pKa -logC)
pH = 1/2(-log(1.4*10^-3)-log(0.01*50))
pH = 1.577
At equivalence point salt is formed hence
pH = 7 + 1/2(pKa + logC)
C = concentration of salt
pH = 7 + 1/2(-log(1.4*10^-3)+log(0.01*50/(50+50))
pH = 7.276